OpenStudy (anonymous):

Find the slope of the line passing through the points (-2, 9) and (5, 3). A. -6/7 B. 4 C. -7/6 D. 1/4

OpenStudy (anonymous):

(3-9)/(5-(-2))=-6/7 =A

OpenStudy (anonymous):

can you help me with this too Find the value of x so that the line passing through the points will have the given slope: (x, 6) and (10, -3) with a slope of -3/2. A. -13 B. 4 C. 13 D. -12

OpenStudy (anonymous):

(-3-6)/(10-x)=-3/2;-9/10-x=-3/2; x=4,B

OpenStudy (anonymous):

thank you